# Classical Simulations of Low Magic Quantum Dynamics

### v1

@@ A.1.2 Pauli Measurements
where $\rho_{S}^{P}$ is a new stabilizer state such that $\rho_{S}^{P},P]=0$ and $\bar{P}$ is a Pauli operator defined analogously to Eq. 30.
@@
Substituting Eqs. 33 and 34 into each term Eq. 32, we see that we arrive at the mapping in Case III.
To evaluate whether $\{P,\mathcal{L}_{\mathrm{LRSD}}\}=0$ in Algorithm 5, and all other commutation relations, we use the symplectic inner product. A generic Pauli operator $P_{i}$ can be written as
@@
where $\oplus$ denotes addition modulo 2. This product equals 0 if $P_{1},P_{2}]=0$ and 1 if $\{P_{1},P_{2}\}=0$.
We now present an approach that leverages the stabilizer structure to efficiently compute the Born probability of projective measurements. The algorithm computes the probability of obtaining the outcome +1+1 when performing a projective measurement of operator $P$ on the general pure state $|\psi\rangle.$ The routine projects the LRSD‐represented state $|\outerproduct{\psi}{\psi}$ onto the +1+1 eigenspace of a Pauli $P$. First, each logical Pauli $\sigma_{l}$ that commutes with $P$ is rewritten as a difference of projectors, allowing the LRSD sum to be expressed entirely in terms of projector products. The trace of this projected operator yields the probability $P_{\uparrow,P}$ of obtaining outcome +1+1. $P_{\uparrow,P}$ can be interpreted as a sum of terms, where each term is a conditional probability of obtaining two consecutive measurements with fixed outcomes. Summing over the terms reconstructs the total measurement probability.
@@
We now present the Algorithm for performing a stabilizer decomposition as in Eq. 31. Firstly, $\rho_{S}^{P}$ is obtained from $\rho_{S}$ by removing a stabilizer generator $\overline{P}$ satisfying $\{\bar{P},P\}=0.$ To see how this decomposition is obtained, we have
@@ A.2 Rényi Entropy of LRSD
We now present algorithms to compute the Rényi entanglement entropies using the representation of the state in Eq. 9. We present an algorithm for computing the reduced density matrix for an arbitrary state, from which we can compute the entanglement. We also present an algorithm which is valid only for $\mathcal{L}_{S},\mathcal{S}]=0.$ These methods allow us to compute the full entanglement spectrum for an arbitrary $n$, where:
@@ A.2.1 General-Purpose Rényi Entanglement Algorithm
We now present a general-purpose algorithm to compute the Rényi entanglement entropies for arbitrary $n$. Given region $A$ with qubits $\{i_{1},\cdots,i_{k}\}$ and region $B$ with qubits $\{i_{1},\cdots,i_{L-k}\},$ the first step is to perform a partial trace of the state in Eq.9 to obtain the reduced density matrix
@@
where $\rho_{S}^{A}$ is the reduced density matrix of $\rho_{S}$ and $\sigma_{\ell^{\prime}}^{A}$ are the logical operators restricted to region $A$. The size of the new logical set $\{\lambda_{\ell^{\prime}}\}$ satisfies $|\{\lambda_{\ell^{\prime}}\}|\leq|\{\lambda_{\ell}\}|.$ Each of the logical operators $\sigma_{\ell}$ and $\rho_{S}^{A}$ can then be converted to the full Hilbert space, represented as $2^{2k}\otimes 2^{2k}$ matrices, respectively. Using matrix operations, the matrix for the full state $\rho_{A}$ can be computed, from which Eq. 40 can be easily evaluated.
@@
We now present a method to perform a partial trace of $\rho$, which is represented using the LRSD in Eq. 9. This method extends the technique in Ref. [35], which shows how to compute the entanglement in stabilizer states. First, we represent the stabilizers of $\rho_{S}$ of 9 in an $r\times L$ matrix, where columns $1$ to $L$ represent qubits $1$ to $L$ and each row represents a different stabilizer. Each entry of the matrix is either the Identity or Pauli operator: II,$X$,$Y$,$Z$. Each logical operator $\sigma_{l}$ is represented analogously. The first step is to bring the tableau of $\rho$ into a reduced row echelon form (RREF), following the procedure in [35]. The general structure of the RREF can be described in a recursive fashion. There are three cases:
@@
The symbols $\vdots$” and $\cdots$” denote a number of repeated rows and columns. This number can be zero. The symbol RREF denotes a possibly empty sub-array that is also in RREF form. The symbol * denotes either a Pauli operator or $I$. Furthermore, $\sigma$, $\sigma_{1}$, and $\sigma_{2}$ are Pauli operators and $\sigma_{1}$ and $\sigma_{2}$ anticommute. We denote the sequences of * operators by $g$, $g_{1}$ and $g_{2}$, respectively. The first column in each of the three cases represents a qubit to be traced over. The stabilizer in RREF’ defines a new stabilizer state $\rho^{\prime}$ which is the reduced density matrix obtained after tracing over the first qubit. In each of the three cases, we can decompose $\rho$ into its action on the qubit to be traced over and the rest of the system. In Case I, recalling that we can substitute the generators $g_{i}$ from the tableau into Eq. 1, we see that:
@@
In Case II, we can also use Eq. 1 to observe that
@@
In Algorithm 5, we present an algorithm for updating the logical set $\mathcal{L}_{S}$ and the destabilizer set $\mathcal{D}$ when performing a measurement of the Pauli operator $P$ in Eq.30. We now describe how to perform a measurement of the Pauli operator $P$, Eq. (30), with the projectors $(I\pm P)/2$, on the stabilizer state $\rho_{S}.$ After measurement, $\pm P$ becomes a stabilizer of $\rho_{S}$, and must therefore be to the stabilizer group. The list of stabilizer generators must then be updated so that they all commute with $P$. Furthermore, the list of destabilizers must be updated to preserve the required commutation relations.
If a single stabilizer $\bar{g}_{i_{1}}$ anticommutes with $P$, we replace $\bar{g}_{i_{1}}$ with $\pm P$ and $\bar{d}_{i_{1}}$ with $\bar{g}_{i_{1}}$. When multiple stabilizers anticommute with $P$, we pair each anticommuting stabilizer and destabilizer into products and perform the updates $\bar{g}_{i_{1}}\mapsto\bar{g}_{i_{1}}\cdot\bar{g}_{i_{k}}$, so the stabilizers now commute with $g_{i_{1}}$. The formation of a list of these products is equivalent to constructing a tree graph, shown in Figure 7, where the nodes represent the anticommuting stabilizers and the edges represent the choices of pairs to multiply.
@@
Step 2: Perform a partial trace over region $B$ for each term in Eq. 51 using Algorithm 4. Then, we rewrite
@@
where each $g_{k}$ is a stabilizer of $\rho_{A}^{(i)}$ and $r_{i}\in\{0,1\}$ defines the phase. Eq. 54 is the probability of successively measuring the $r_{i}$ stabilizers of $\rho_{A}^{(i)}$ with the outcomes $\pm 1$ encoded via the $\{r_{k}\}.$ If $\rho_{A}^{(i)}$ and $\rho_{A}^{(j)}$ are pure states, then $\rho_{A}^{(i)}=|\psi_{i}\rangle\langle\psi_{i}|$ and $\rho_{A}^{(j)}=|\psi_{j}\rangle\langle\psi_{j}|.$ Then, $\mathrm{tr}[\rho_{A}^{(i)}\rho_{A}^{(j)}]=|\langle\psi_{i}|\psi_{j}\rangle|^{2}$ can be computed using the inner product formula in Ref.[10]. Namely, enumerate the stabilizer groups $\mathcal{S}_{\psi_{i}}$ and $\mathcal{S}_{\psi_{j}}$ for $|\psi_{i}\rangle$ and $|\psi_{j}\rangle.$ Let
@@
Case II: For arbitrary $n$, each term in Eq. 52 can be represented as a matrix on the full Hilbert space, and thus $\rho_{A}$ can be constructed explicitly. The entanglement can then be computed directly from $\rho_{A}$.
@@ A.3.2 Bell Sampling Algorithm
We now introduce a Monte Carlo sampling algorithm that efficiently produces Bell samples for a specific subclass of states that include, but are not limited to, stabilizer states. The sampler only visits Bell samples, $\sigma_{\vec{r}}$, which have nonzero probabilities $p(\vec{r})>0$, where $p(\vec{r})$ is defined in Eq. 58. This algorithm is adapted from Ref. [37]. The non-stabilizer state must be able to be written as $|\psi\rangle=\mathcal{T}\mathcal{C}|\psi_{o}\rangle$, where $|\psi_{o}\rangle$ is a stabilizer state, $\mathcal{C}$ is a Clifford circuit, and $\mathcal{T}$ is a layer of $T$-gates. The non-Clifford layer factorizes as
@@
We now consider how the density matrix $\rho_{S}$, for the initial state $\mathcal{C}|\psi_{o}\rangle$, evolves under the $T$-gate layer in Eq. 59. The final state $\rho$ can be represented as a product of a mixed stabilizer state and a product over $k$ terms, each of which is written as a sum over three projectors:
@@ A.3.3 Magic Computation
We now present an algorithm for Bell sampling from the state in Eq. 61. Recall from Section A.3.1 that a Bell sample $\sigma_{\vec{r}}$ is obtained from $|\psi\rangle\otimes|\psi\rangle$ with probability $p(\vec{r})$. This is equivalent to measuring each qubit pair in the Bell basis and obtaining the outcome sequence $\vec{r}$ with Born probability $p(\vec{r})$. We now define $\sigma_{\vec{r}_{o}}$ such that
@@
Step 2: Compute $\sigma_{\vec{r}_{o}}$, as defined in Eq. 65,
@@
To test the accuracy of Algorithm 10, we fix the location of the unitary gates and measurement locations for many circuits. We sample all $4^{L}$ possible Pauli operators on $L$ qubits. We then compute $\mathcal{M}_{\mathrm{exact}}$ using Eq. (64) with these samples. We also obtain a numerical estimate $\mathcal{M}_{\mathrm{numerical}}$ for the same set of circuits, using Algorithm 10 with $\tau$ samples. Then, define $|\delta\mathcal{M}|=\mathcal{M}_{\mathrm{numerical}}-\mathcal{M}_{\mathrm{exact}}.$ The absolute value $|\delta\mathcal{M}|$ will then be positive and decrease to 0 with increasing $\tau$, as shown in Figure 9.
@@
We now demonstrate our algorithm on the single-pair all-to-all model with $Z$-basis measurements. From the discussion at the beginning of Section IV.1, we observe that the state can be factorized with the $T$-gate layer in Eq. 59 acting at the final timestep.
@@
Figure 9: (a) Absolute value of the difference in the exact stabilizer nullity, defined in Eq. (64), and the numerical estimate, $|\absolutevalue{\delta\overline{\mathcal{M}}}$ as a function of the sampling time $\tau$. $\tau$ is defined in Algorithm 10. $\overline{\phantom{X}}$ denotes averaging over circuits. The exact $\mathcal{M}$ for each circuit is obtained by sampling all possible Pauli operators to compute Eq. (64), and the numerical estimate is obtained using $\tau$ sampling steps of Algorithm 10. Data is obtained in the single-pair all-to-all model with $Z$-basis measurements, system size $L=48$, and measurement probabilities $p_{m}\in\{0.0,\ 0.2,\ 0.4,\ 0.6,\ 0.8\}$. We average over $\mathcal{O}(10^{3})$ iterations to obtain each curve. For each fixed $p_{m}$ and circuit, the locations of unitary gates and measurements are the same for the numerical and exact estimates of $\mathcal{M}$.
@@
Figure 10 illustrates the convergence behavior of the stabilizer nullity $\mathcal{M}$ and the number of distinct Bell samples $\mathcal{N}_{s}$ as a function of the number of samples $\tau$, for different measurement rates $p_{m}.$ In the top row, we observe that $\mathcal{M}$ exhibits rapid growth followed by saturation for all values of $p_{m}.$ The saturation value decreases as $p_{m}$ increases, reflecting expected suppression of non-stabilizerness by projective measurements. Additionally, saturation occurs earlier for larger $p_{m}$, which correlates with the reduced complexity of the state. The bottom row, $\mathcal{N}_{s}$, the number of distinct Bell samples obtained during sampling, is independent of the measurement rate $p_{m}.$ It grows at a rate of approximately one per sampling step, indicating that it is rare for the same Pauli to be sampled twice in the same Monte Carlo run.
@@
Figure 10: The top row shows the stabilizer nullity $\mathcal{M}$, defined in Eq. (64), as a function of the sampling time $\tau$ for fixed measurement probabilities $p_{m}\in\{0.0,0.2,0.4,0.6,0.8\}$. $\overline{\phantom{X}}$ denotes averaging over circuits. We use Algorithm 10 to compute $\mathcal{M}$, where $\tau$ is defined in 10. Data is obtained from the single-pair all-to-all model with $X$-basis measurements. Each curve represents an average over $\mathcal{O}(10^{3})$ circuit realizations. Results are shown for system sizes: (a) $L=16$, (b) $L=32$, (c) $L=48$, and (d) $L=64$. The bottom row, panels (e-h), shows the number of distinct Bell samples $\mathcal{N}_{s}$, defined in Step 1 of Algorithm 10, as a function of the sampling time $\tau$. Data was obtained using the same set of circuits as in the top row, and averaged over $\mathcal{O}(\mathrm{10}^{3})$ trajectories.

### v2

@@ A.1.2 Pauli Measurements
where $\rho_{S}^{P}$ is a new stabilizer state such that $\rho_{S}^{P},P]=0$ and $\bar{P}$ is a Pauli operator defined analogously to Eq. (28).
@@
Substituting Eqs. (31) and  (32) into each term of Eq. (30), we see that we arrive at the mapping in Case III. To evaluate whether $\{P,\mathcal{L}_{\mathrm{LRSD}}\}=0$ in Algorithm 5, and all other commutation relations, we use the symplectic inner product. A generic Pauli operator $P_{i}$ can be written as
@@
where $\oplus$ denotes addition modulo 2. This product equals 0 if $P_{1},P_{2}]=0$ and 1 if $\{P_{1},P_{2}\}=0$. We now present an approach that leverages the stabilizer structure to efficiently compute the Born probability of projective measurements. The algorithm computes the probability of obtaining the outcome +1+1 when performing a projective measurement of operator $P$ on the general pure state $|\psi\rangle.$ The routine projects the LRSD‐represented state $|\outerproduct{\psi}{\psi}$ onto the +1+1 eigenspace of a Pauli $P$. First, each logical Pauli $\sigma_{l}$ that commutes with $P$ is rewritten as a difference of projectors, allowing the LRSD sum to be expressed entirely in terms of projector products. The trace of this projected operator yields the probability $P_{\uparrow,P}$ of obtaining outcome +1+1. $P_{\uparrow,P}$ can be interpreted as a sum of terms, where each term is a conditional probability of obtaining two consecutive measurements with fixed outcomes. Summing over the terms reconstructs the total measurement probability.
@@
We now present the Algorithm for performing a stabilizer decomposition as in Eq. (29). Firstly, $\rho_{S}^{P}$ is obtained from $\rho_{S}$ by removing a stabilizer generator $\overline{P}$ satisfying $\{\bar{P},P\}=0.$ To see how this decomposition is obtained, we have
@@ A.2 Rényi Entropy of LRSD
We now present algorithms to compute the Rényi entanglement entropies using the representation of the state in Eq. (4). We present an algorithm for computing the reduced density matrix for an arbitrary state, from which we can compute the entanglement. We also present an algorithm which is valid only for $\mathcal{L}_{S},\mathcal{S}]=0.$ These methods allow us to compute the full entanglement spectrum for an arbitrary $n$, where:
@@ A.2.1 General-Purpose Rényi Entanglement Algorithm
We now present a general-purpose algorithm to compute the Rényi entanglement entropies for arbitrary $n$. Given region $A$ with qubits $\{i_{1},\cdots,i_{k}\}$ and region $B$ with qubits $\{i_{1},\cdots,i_{L-k}\},$ the first step is to perform a partial trace of the state in Eq. (4) to obtain the reduced density matrix
@@
where $\rho_{S}^{A}$ is the reduced density matrix of $\rho_{S}$ and $\sigma_{\ell^{\prime}}^{A}$ are the logical operators restricted to region $A$. The size of the new logical set $\{\lambda_{\ell^{\prime}}\}$ satisfies $|\{\lambda_{\ell^{\prime}}\}|\leq|\{\lambda_{\ell}\}|.$ Each of the logical operators $\sigma_{\ell}$ and $\rho_{S}^{A}$ can then be converted to the full Hilbert space, represented as $2^{2k}\otimes 2^{2k}$ matrices, respectively. Using matrix operations, the matrix for the full state $\rho_{A}$ can be computed, from which Eq.  (38) can be easily evaluated.
@@
We now present a method to perform a partial trace of $\rho$, which is represented using the LRSD in Eq. (4). This method extends the technique in Ref. [43], which shows how to compute the entanglement in stabilizer states. First, we represent the stabilizers of $\rho_{S}$ of Eq. (8) in an $r\times L$ matrix, where columns $1$ to $L$ represent qubits $1$ to $L$ and each row represents a different stabilizer. Each entry of the matrix is either the Identity or Pauli operator: II,$X$,$Y$,$Z$. Each logical operator $\sigma_{l}$ is represented analogously. The first step is to bring the tableau of $\rho$ into a reduced row echelon form (RREF), following the procedure in [43]. The general structure of the RREF can be described in a recursive fashion. There are three cases:
@@
The symbols $\vdots$” and $\cdots$” denote a number of repeated rows and columns. This number can be zero. The symbol RREF denotes a possibly empty sub-array that is also in RREF form. The symbol * denotes either a Pauli operator or $I$. Furthermore, $\sigma$, $\sigma_{1}$, and $\sigma_{2}$ are Pauli operators and $\sigma_{1}$ and $\sigma_{2}$ anticommute. We denote the sequences of * operators by $g$, $g_{1}$ and $g_{2}$, respectively. The first column in each of the three cases represents a qubit to be traced over. The stabilizer in RREF’ defines a new stabilizer state $\rho^{\prime}$ which is the reduced density matrix obtained after tracing over the first qubit. In each of the three cases, we can decompose $\rho$ into its action on the qubit to be traced over and the rest of the system. In Case I, recalling that we can substitute the generators $g_{i}$ from the tableau into Eq. (1), we see that:
@@
In Case II, we can also use Eq. (1) to observe that
@@
In Algorithm 5, we present an algorithm for updating the logical set $\mathcal{L}_{S}$ and the destabilizer set $\mathcal{D}$ when performing a measurement of the Pauli operator $P$ in Eq. (28). We now describe how to perform a measurement of the Pauli operator $P$, Eq. (28), with the projectors $(I\pm P)/2$, on the stabilizer state $\rho_{S}.$ After measurement, $\pm P$ becomes a stabilizer of $\rho_{S}$, and must therefore be to the stabilizer group. The list of stabilizer generators must then be updated so that they all commute with $P$. Furthermore, the list of destabilizers must be updated to preserve the required commutation relations.
If a single stabilizer $\bar{g}_{i_{1}}$ anticommutes with $P$, we replace $\bar{g}_{i_{1}}$ with $\pm P$ and $\bar{d}_{i_{1}}$ with $\bar{g}_{i_{1}}$. When multiple stabilizers anticommute with $P$, we pair each anticommuting stabilizer and destabilizer into products and perform the updates $\bar{g}_{i_{1}}\mapsto\bar{g}_{i_{1}}\cdot\bar{g}_{i_{k}}$, so the stabilizers now commute with $g_{i_{1}}$. The formation of a list of these products is equivalent to constructing a tree graph, shown in Fig. 7, where the nodes represent the anticommuting stabilizers and the edges represent the choices of pairs to multiply.
@@
Step 2: Perform a partial trace over region $B$ for each term in Eq. (49) using Algorithm 4. Then, we rewrite
@@
where each $g_{k}$ is a stabilizer of $\rho_{A}^{(i)}$ and $r_{i}\in\{0,1\}$ defines the phase. Eq. (52) is the probability of successively measuring the $r_{i}$ stabilizers of $\rho_{A}^{(i)}$ with the outcomes $\pm 1$ encoded via the $\{r_{k}\}.$ If $\rho_{A}^{(i)}$ and $\rho_{A}^{(j)}$ are pure states, then $\rho_{A}^{(i)}=|\psi_{i}\rangle\langle\psi_{i}|$ and $\rho_{A}^{(j)}=|\psi_{j}\rangle\langle\psi_{j}|.$ Then, $\mathrm{tr}[\rho_{A}^{(i)}\rho_{A}^{(j)}]=|\langle\psi_{i}|\psi_{j}\rangle|^{2}$ can be computed using the inner product formula in Ref.[10]. Namely, enumerate the stabilizer groups $\mathcal{S}_{\psi_{i}}$ and $\mathcal{S}_{\psi_{j}}$ for $|\psi_{i}\rangle$ and $|\psi_{j}\rangle.$ Let
@@
Case II: For arbitrary $n$, each term in Eq. (50) can be represented as a matrix on the full Hilbert space, and thus $\rho_{A}$ can be constructed explicitly. The entanglement can then be computed directly from $\rho_{A}$.
@@ A.3.2 Bell Sampling Algorithm
We now introduce a Monte Carlo sampling algorithm that efficiently produces Bell samples for a specific subclass of states that include, but are not limited to, stabilizer states. The sampler only visits Bell samples, $\sigma_{\vec{r}}$, which have nonzero probabilities $p(\vec{r})>0$, where $p(\vec{r})$ is defined in Eq.  (56). This algorithm is adapted from Ref. [45]. The non-stabilizer state must be able to be written as $|\psi\rangle=\mathcal{T}\mathcal{C}|\psi_{o}\rangle$, where $|\psi_{o}\rangle$ is a stabilizer state, $\mathcal{C}$ is a Clifford circuit, and $\mathcal{T}$ is a layer of $T$-gates. The non-Clifford layer factorizes as
@@
We now consider how the density matrix $\rho_{S}$, for the initial state $\mathcal{C}|\psi_{o}\rangle$, evolves under the $T$-gate layer in Eq. (57). The final state $\rho$ can be represented as a product of a mixed stabilizer state and a product over $k$ terms, each of which is written as a sum over three projectors:
@@ A.3.3 Magic Computation
We now present an algorithm for Bell sampling from the state in Eq. (59). Recall from Section A.3.1 that a Bell sample $\sigma_{\vec{r}}$ is obtained from $|\psi\rangle\otimes|\psi\rangle$ with probability $p(\vec{r})$. This is equivalent to measuring each qubit pair in the Bell basis and obtaining the outcome sequence $\vec{r}$ with Born probability $p(\vec{r})$. We now define $\sigma_{\vec{r}_{o}}$ such that
@@
Step 2: Compute $\sigma_{\vec{r}_{o}}$, as defined in Eq.  (63),
@@
To test the accuracy of Algorithm 10, we fix the location of the unitary gates and measurement locations for many circuits. We sample all $4^{L}$ possible Pauli operators on $L$ qubits. We then compute $\mathcal{M}_{\mathrm{exact}}$ using Eq. (62) with these samples. We also obtain a numerical estimate $\mathcal{M}_{\mathrm{numerical}}$ for the same set of circuits, using Algorithm 10 with $\tau$ samples. Then, define $|\delta\mathcal{M}|=\mathcal{M}_{\mathrm{numerical}}-\mathcal{M}_{\mathrm{exact}}.$ The absolute value $|\delta\mathcal{M}|$ will then be positive and decrease to 0 with increasing $\tau$, as shown in Fig. 9.
@@
We now demonstrate our algorithm on the single-pair all-to-all model with $Z$-basis measurements. From the discussion at the beginning of Section IV.1, we observe that the state can be factorized with the $T$-gate layer in Eq.  (57) acting at the final timestep.
@@
Figure 9: Absolute value of the difference in the exact stabilizer nullity, defined in Eq. (62), and the numerical estimate, $|\absolutevalue{\delta\overline{\mathcal{M}}}$ as a function of the sampling time $\tau$. $\tau$ is defined in Algorithm 10. $\overline{\cdots}$ denotes averaging over circuit trajectories. The exact $\mathcal{M}$ for each circuit is obtained by sampling all possible Pauli operators to compute Eq. (62), and the numerical estimate is obtained using $\tau$ sampling steps of Algorithm 10. Data is obtained in the single-pair all-to-all model with $Z$-basis measurements, system size $L=48$, and measurement probabilities $p_{m}\in\{0.0,\ 0.2,\ 0.4,\ 0.6,\ 0.8\}$. We average over $\mathcal{O}(10^{3})$ iterations to obtain each curve. For each fixed $p_{m}$ and circuit, the locations of unitary gates and measurements are the same for the numerical and exact estimates of $\mathcal{M}$.
@@
Fig. 10 illustrates the convergence behavior of the stabilizer nullity $\mathcal{M}$ and the number of distinct Bell samples $\mathcal{N}_{s}$ as a function of the number of samples $\tau$, for different measurement rates $p_{m}.$ In the top row, we observe that $\mathcal{M}$ exhibits rapid growth followed by saturation for all values of $p_{m}.$ The saturation value decreases as $p_{m}$ increases, reflecting expected suppression of non-stabilizerness by projective measurements. Additionally, saturation occurs earlier for larger $p_{m}$, which correlates with the reduced complexity of the state. The bottom row, $\mathcal{N}_{s}$, the number of distinct Bell samples obtained during sampling, is independent of the measurement rate $p_{m}.$ It grows at a rate of approximately one per sampling step, indicating that it is rare for the same Pauli to be sampled twice in the same Monte Carlo run.
@@
Figure 10: The top row shows the stabilizer nullity $\mathcal{M}$, defined in Eq. (62), as a function of the sampling time $\tau$ for fixed measurement probabilities $p_{m}\in\{0.0,0.2,0.4,0.6,0.8\}$. $\overline{\cdots}$ denotes averaging over circuit trajectories. We use Algorithm 10 to compute $\mathcal{M}$, where $\tau$ is defined in 10. Data is obtained from the single-pair all-to-all model with $X$-basis measurements. Each curve represents an average over $\mathcal{O}(10^{3})$ circuit realizations. Results are shown for system sizes: (a) $L=16$, (b) $L=32$, (c) $L=48$, and (d) $L=64$. The bottom row, panels (e-h), shows the number of distinct Bell samples $\mathcal{N}_{s}$, defined in Step 1 of Algorithm 10, as a function of the sampling time $\tau$. Data was obtained using the same set of circuits as in the top row, and averaged over $\mathcal{O}(\mathrm{10}^{3})$ trajectories.